# Understanding trade-offs in classifier bias with quality-diversity optimization: an application to talent management

### v1

@@ 2.3 Quality-diversity search
Map-Elites, like other forms of evolutionary computation, is a stochastic search/optimization algorithm [8]. This means that it is at efficiency disadvantage compared to methods for model optimization tailored to specific learner representations, such as gradient descent (for neural network training) or information gain maximization (for decision tree learning). However, it has the advantage that stochastic search can be applied to optimize anything that has parameters and an objective function. Common learner representations such as neural networks and decision trees can be evolved with good results, though this process is often less computationally efficient than using model-specific learning methods.
@@ 3 Methodology
Next, the weights of a set of neural networks with the defined architecture are evolved, and their fitness values are calculated as the accuracy of the network prediction against the real outcome for the whole dataset.
The above-mentioned accuracy is used as the output value to be maximized, while the ratio of positive classification between the two groups present in each protected attribute are used as descriptors of the model (for example the ratio of female positive prediction rate vs male positive prediction rate for gender attribute, or the ratio of young positive rate vs old positive rate for age.)
@@
The used code relies on the pyribs [30] library.
@@ 3.1 Data
The Promotion Dataset, is an open source dataset consisting of individuals that were either promoted or not promoted and a set of features including department in the organization, region, education, gender, training, age, performance, and tenure. Categorical variables were encoded using binary dummy variables.
We used sampling to generate different scenarios from the data, to evaluate how the presence or absence of bias in the data impacts the results. Random sampling was used to obtain 4 datasets with a defined bias. In all cases the total percentage of promoted individuals is set close to 50%. Given the age distribution in the data, a 34 years old threshold was used to define the young vs. old groups.
Tables 2 to 5 show the descriptions of the data used, including the number of cases for each group, the proportion of cases for each group within its category (gender and age), and the promotion rate for the group in total. Additionally we showed how promotion rate changes for each group in a category considering the distribution in the other category.
Adult Dataset, predicts annual income exceeding of \$50K/yr using census data. Features include age, gender, race, education, occupation, among others. This dataset has bias for both protected attributes. Random sampling was used to obtain a stratified sample with similar distribution to the original dataset and a unbiased sample.
Tables 6 to 7 show the sample details for Adult dataset.
@@
    Promotion Unbiased dataset, with 50% promotion rate for all the categories, Table 2.
    Table 2: Unbiased dataset
    n Proportion Promotion
    rate
    All 5728 1.00 0.5
    Male 2864 0.5 0.5
    Female 2864 0.5 0.5
    Young 2864 0.5 0.5
    Old 2864 0.5 0.5
    Promotion Young Old
    rate
    Male 0.5 0.5
    Female 0.5 0.5
@@
    Promotion dataset with same proportion for male, female, young, old; with higher promotion rate for male in both age groups, Table 3
    Table 3: Male biased dataset
    n Proportion Promotion
    rate
    All 4384 1.00 0.5
    Male 2192 0.5 0.65
    Female 2192 0.5 0.35
    Young 2192 0.5 0.5
    Old 2192 0.5 0.5
    Promotion Young Old
    rate
    Male 0.65 0.65
    Female 0.35 0.35
@@
    Promotion dataset with higher proportion for male; similar proportion for young, old; with higher promotion rate for male in both age groups. The sample size for male is bigger than for female, Table 4
    Table 4: Male biased dataset, higher male sample
    n Proportion Promotion
    rate
    All 6576 1.00 0.56
    Male 4384 0.67 0.67
    Female 2192 0.33 0.35
    Young 3288 0.5 0.56
    Old 3288 0.5 0.56
    Promotion Young Old
    rate
    Male 0.67 0.67
    Female 0.33 0.33
@@
    Promotion dataset with same proportion for male, female; similar proportion for young, old; with higher promotion rate for male if young and higher promotion rate for female if old, Table 5
    Table 5: Cross biased dataset, bias for young male and old female
    n Proportion Promotion
    rate
    All 4508 1.00 0.5
    Male 2254 0.5 0.5
    Female 2254 0.5 0.5
    Young 2254 0.5 0.5
    Old 2254 0.5 0.5
    Promotion Young Old
    rate
    Male 0.33 0.67
    Female 0.67 0.33
@@
    Adult unbiased dataset, Table 6
    Table 6: Adult Unbiased Sample
    n Proportion Promotion
    rate
    All 1680 1.00 0.5
    Male 420 0.5 0.5
    Female 420 0.5 0.5
    White 420 0.5 0.5
    Other Race 420 0.5 0.5
    Promotion White Other
    rate
    Male 0.5 0.5
    Female 0.5 0.5
@@
    Adult stratified sample, Table 7
    Table 7: Adult Stratified Sample
    n Proportion Promotion
    rate
    All 13565 1.00 0.25
    Male 9224 0.68 0.31
    Female 4341 0.32 0.11
    White 11666 0.86 0.26
    Other Race 1899 0.14 0.16
    Promotion White Other
    rate
    Male 0.3239 0.2243
    Female 0.1223 0.0759
@@ 4 Results
In this sample (see Table 2), we have the same number of promoted individuals in all the groups: male, female, old and young individuals, making the promotion rate 50% for all cases.
@@ 4.2 Male biased sample (Figure 3, Table 9)
While the number of male and female, and the number of young and old individuals is equally split, the proportion of males in the sample that are promoted is 65% and 35% for females. Regarding age, both groups have a similar promotion rate of 50% (Table 3.)
@@ 4.3 Male biased sample, male oversampling (Figure 4, Table 10)
For this case, we wanted to observe the effect of having a higher promotion rate for males and a higher number of male cases represented in the dataset (Table 4), a condition found in many real life situations. Both age groups have a similar promotion rate.
@@
In this experiment we used a dataset that doesn’t have bias considering each dimension independently (male vs. female or young vs old). But, when we combined the dimensions, there is bias: individuals that are male and young have a higher promotion rate, so do individuals that are female and young (Table 5)
@@
In this experiment we used a unbiased sample from the Adult dataset (Table 6)
@@
In this experiment we used a stratified sample from the Adult dataset that corresponds to the distribution of positive outcomes for both groups in the protected attributes(Table 7)
@@ 5 Discussion
While there are approaches that seek to alleviate biases rather than characterize them, they all have various shortcoming. For example, dominant classes could be undersampled, but that means that not all of the dataset will be used, likely impacting classifier performance and generalization; it also requires that an analysis of bias is done in the first place to understand which instances to undersample.
@@ 6 Conclusion and Future Work
We showed that CMA-ME can generate multiple models of varying bias, thereby allowing us to visualize the distribution of the accuracy of the models and its relation with the learned bias. In other words, we can understand the bias-performance trade-off. The result of the method is a set of trained models, where the user can choose any of them for deployment. The accuracy of the best-classifying model is comparable, though slightly inferior to, the output of state of the art convenational machine learning methods. The method is fast and would scale to networks of at least tens of thousands of parameters.

### v2

@@ 2.3 Quality-diversity search
Map-Elites, like other forms of evolutionary computation, is a stochastic search/optimization algorithm [8]. This means that it is at efficiency disadvantage compared to methods for model optimization tailored to specific learner representations, such as gradient descent (for neural network training) or information gain maximization (for decision tree learning). However, it has the advantage that stochastic search can be applied to optimize anything that has parameters and an objective function. Common learner representations such as neural networks and decision trees can be evolved with good results [33], though this process is often less computationally efficient than using model-specific learning methods.
@@ 3 Methodology
Next, the weights of a set of neural networks with the defined architecture are evolved, and their fitness values are calculated as the accuracy of the network forward pass prediction against the real outcome for the whole dataset.
The above-mentioned accuracy is used as the output value to be maximized, while the ratio of positive classification between the two groups present in each protected attribute are used as descriptors of the model (for example the ratio of female positive prediction rate vs male positive prediction rate for gender attribute, or the ratio of young positive rate vs old positive rate for age.) See Algorithm 2.
Algorithm 2 Fitness Function
1:  Input: weights, nn_architecture, cases, cases_xa, cases_xb, cases_ya, cases_yb, labels.
2:  Output: accuracy, ratio_x, ratio_y.
3:  Step 1: Compute Overall Accuracy
4:  Perform a forward pass on cases with weights and nn_architecture.
5:  Compute accuracy:
accuracycorrect predictionstotal $\texttt{accuracy}\leftarrow\frac{\texttt{correct predictions}}{\texttt{total cases}}$
6:  Step 2: Compute Group Ratios
7:  For each group (cases_xa, cases_xb, cases_ya, cases_yb), compute mean predictions:
[…]
8:  Compute fairness ratios:
[…]
9:  Step 3: Return Fitness Results
10:  Output: (accuracy, (ratio_x, ratio_y))
@@
NN Activation function Leaky Relu Leaky Relu
The used code relies on the pyribs [31] library. Details about the CMA-ME parameters and code can be found at https://github.com/catajara/Classif-Bias-Tradeoffs-QD
@@ 3.1 Data
The Promotion Dataset, is an open source dataset consisting of individuals that were either promoted or not promoted and a set of features including department in the organization, region, education, gender, training, age, performance, and tenure. Categorical variables were encoded using binary dummy variables. We used sampling to generate different scenarios from the data, to evaluate how the presence or absence of bias in the data impacts the results. Random sampling was used to obtain 4 datasets with a defined bias. In all cases the total percentage of promoted individuals is set close to 50%. Given the age distribution in the data, a 34 years old threshold was used to define the young vs. old groups. Rows 1 to 4 in Table 2 show the data used, including the number of cases for each group and Row 1 to 4 in Table 3 show the promotion rate for the group:
@@
    Promotion Unbiased dataset, with 50% promotion rate for all the categories.
@@
    Promotion dataset with same proportion for male and female, and for young and old; with higher promotion rate for male cases in both age groups.
@@
    Promotion dataset with higher sample size for male cases, and similar proportion for young and old groups; with higher promotion rate for male in both age groups.
@@
    Promotion dataset with same proportion for male and female cases, and similar proportion for young and old; with higher promotion rate for male and young cases (67% compared with 33% for male and old), and higher promotion rate for female and old (67%) vs. female and young cases (33%).
Adult Dataset, predicts annual income exceeding of \$50K/yr using census data. Features include age, gender, race, education, occupation, among others. This dataset has bias for both protected attributes (gender and race in this case). Random sampling was used to obtain a stratified sample with similar distribution to the original dataset and a unbiased sample. Rows 5 and 6 in Tables 2 and 3 show the sample details for Adult dataset:
@@
    Adult unbiased dataset , with 50% promotion rate for all the categories.
@@
    Adult stratified sample, with higher number cases for male and for white, and higher positive rate for male and for white.
Table 2: Dataset Information
No. Dataset Total (n) Male (n) Female (n) Young/ Old
White (n) Other (n)
1 Unbiased 5728 2864 2864 2864 2864
2 Male Biased 4384 2192 2192 2192 2192
3 Higher Male Sample 6576 4384 2192 3288 3288
4 Cross Biased 4508 2254 2254 2254 2254
5 Adult Unbiased 1680 840 840 840 840
6 Adult Stratified 13565 9224 4341 11666 1899
Table 3: Positive Rate by Subgroup
No. Dataset All Male Female Young/White Old/Other
1 Unbiased 0.50 0.50 0.50 0.50 0.50
2 Male Biased 0.50 0.65 0.35 0.50 0.50
3 Higher Male Sample 0.56 0.67 0.35 0.56 0.56
4 Cross Biased 0.50 0.50 0.50 0.50 0.50
5 Adult Unbiased 0.50 0.50 0.50 0.50 0.50
6 Adult Stratified 0.25 0.31 0.11 0.26 0.16
@@ 4 Results
In this sample (see Row 1 Table 3), we have the same number of promoted individuals in all the groups: male, female, old and young individuals, making the promotion rate 50% for all cases.
@@ 4.2 Male biased sample (Figure 3, Table 5)
While the number of male and female, and the number of young and old individuals is equally split, the proportion of males in the sample that are promoted is 65% and 35% for females. Regarding age, both groups have a similar promotion rate of 50% (Row 2 Table 3.)
@@ 4.3 Male biased sample, male oversampling (Figure 4, Table 6)
For this case, we wanted to observe the effect of having a higher promotion rate for males and a higher number of male cases represented in the dataset (Row 2 Tables 2 and 3), a condition found in many real life situations. Both age groups have a similar promotion rate.
@@
In this experiment we used a dataset that doesn’t have bias considering each dimension independently (male vs. female or young vs old). But, when we combined the dimensions, there is bias: individuals that are male and old have a higher promotion rate, so do individuals that are female and young (Row 4 Table 2)
@@
In this experiment we used a unbiased sample from the Adult dataset (Row 5 Table 3)
@@
In this experiment we used a stratified sample from the Adult dataset that corresponds to the distribution of positive outcomes for both groups in the protected attributes (gender and race). See Row 6 Tables 2 and 3).
@@ 5 Discussion
While there are approaches that seek to alleviate biases rather than characterize them, they all have various shortcomings. For example, dominant classes could be undersampled, but that means that not all of the dataset will be used, likely impacting classifier performance and generalization; it also requires that an analysis of bias is done in the first place to understand which instances to undersample.
@@ 6 Conclusion and Future Work
We showed that CMA-ME can generate multiple models of varying bias, thereby allowing us to visualize the distribution of the accuracy of the models and its relation with the learned bias. In other words, we can understand the bias-performance trade-off. The result of the method is a set of trained models, where the user can choose any of them for deployment. The accuracy of the best-classifying model is comparable, though slightly inferior to, the output of state of the art conventional machine learning methods. The method is fast and would scale to networks of at least tens of thousands of parameters [30].